the motion of a particle is given by x(t)=(25cm)cos(13t) where t is in s

The motion of a particle is given by x(t)=(25cm)cos(13t), where t is in s.

At what time is the kinetic energy equal to twice the potential energy for the first time?
Answer #1

KE = 2PE


PE = (1/2) k x^2

KE = (1/2) kA^2 - (1/2)kx^2

(1/2) kA^2 - (1/2)kx^2 = 2 * (1/2) kx^2


A^2 = 3x^2

(25)^2 = 3 * [(25cm) cos(13t)]^2

cos^2 13t = 1/3

cos 13 t = 1/?3

13 t = 0.9553

t = 0.07 s.